INFRARED SPECTRA OF THE 2-CHLOROPROPYL RADICAL IN SOLID PARA-HYDROGEN

\underline{JAY C. AMICANGELO}, School of Science, Penn State Erie, Erie, PA 16563; BARBARA GOLEC and YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

The reaction of chlorine atoms with propylene and one of its deuterium isotopomers in solid para-hydrogen (p-H_{2}) matrices at 3.2 K has been studied using infrared spectroscopy. Irradiation at 365 nm of a co-deposited mixture of Cl_{2}, C_{3}H_{6}, and p-H_{2} at 3.2 K produces a series of new lines in the infrared spectrum. Several of the new lines are readily assigned to the gauche and trans conformers of 1,2-dichloropropane (CH_{3}CHClCH_{2}Cl) resulting from the addition of two Cl atoms to C_{3}H_{6}. Weak lines observed at 802 and 975 cm^{−1} and at 813 and 981 cm^{−1} that become more prominent upon secondary irradiation at 254 and 214 nm are assigned to the allyl radical (C_{3}H_{5}) and an HCl-allyl radical complex (HCl-C_{3}H_{5}), respectively^{a}. Of the remaining lines, a strong line at 650 cm^{−1} and weaker lines at 532, 1008, 1133, 1150, 1215 and 1382 cm^{−1} are concluded to be due to a single carrier based on their behavior upon subsequent annealing to 4.5 K and irradiation at 254 and 214 nm. When the positions and intensities of these lines are compared to the MP2/aug-cc-pVDZ predicted vibrational spectra of the possible species that could result from the addition and abstraction reactions of one Cl atom with C_{3}H_{6}^{a}, the best agreement is found with the 2-chloropropyl radical (CH_{3}CHClCH_{2}·). Isotopic experiments were performed with 3,3,3-C_{3}H_{3}D_{3} and the corresponding infrared peaks due to the deuterium isotopomer of this radical (CD_{3}CHClCH_{2}·) have also been observed. A final set of experiments were performed following irradiation of the Cl_{2}/C_{3}H_{6}/p-H_{2} mixture at 365 nm, in which the matrix was irradiated with filtered infrared light from a globar source, which has been shown to induce a reaction between isolated Cl atoms and matrix H_{2} to produce HCl and H atoms^{b}. In our experiments, the major products observed after infrared irradiation are HCl, 2-chloropropane (CH_{3}CHClCH_{3}) and the isopropyl radical (CH_{3}CH·CH_{3}) and the possible mechanisms of formation of these species will be discussed.

aP. Brana and J. A. Sordo, J. Comput. Chem. \underline{24}, 2044 (2003)

bP. L. Raston and D. T. Anderson, Phys. Chem. Chem. Phys. \underline{8}, 3124 (2006)